OpenStudy (anonymous):

Someone PLEASE help? I don't get these problem! 1.) find the standard form or the equation of the ellipse with vertices (±9,0) and eccentricity e=1/3. 2.) Find the standard form of the equation of the hyperbola with the given characteristics. vertices: (2, - 7), (2, - 1), foci: (2, - 9), (2, 1) 3.) Find the standard form of the equation of the ellipse with the given characteristics. center: (-1, 7) a = 3c foci: (- 2, 7), (0, 7

OpenStudy (sumi29):

Well does your course cover conics? If it does, then you should be familiar with the standard equations of ellipse and hyperbola, and then its a simple matter of substitution

OpenStudy (anonymous):

center is half way beteen \((-9,0)\) and \((9,0)\) so the center is \((0,0)\) which makes things easier

OpenStudy (anonymous):

equation will look like \(\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\) and what you need now is \(a\) and \(b\)

OpenStudy (anonymous):

so this isnt an x-h/a+y-k/b?

OpenStudy (anonymous):

oh wait now i see. sorry

OpenStudy (anonymous):

no since \(h=k=0\) in this case

OpenStudy (anonymous):

now you have \(a=9\) because the vertices are \((-9,0)\) and \((9,0)\)

OpenStudy (anonymous):

\[\frac{x^2}{81}+\frac{y^2}{b^2}=1\]

OpenStudy (anonymous):

now you need \(b\)

OpenStudy (anonymous):

is b 9?

OpenStudy (anonymous):

idk we have to find it, but i doubt it

OpenStudy (anonymous):

ok how do you find b?

OpenStudy (anonymous):

you have the eccentricity as \(\frac{1}{3}\) right?

OpenStudy (anonymous):

yeah its 1/3

OpenStudy (anonymous):

and we also know \(a=9\) in general the eccentricity is \(\frac{c}{a}\) where \(c=\sqrt{a^2-b^2}\) or \(c^2=a^2-b^2\)

OpenStudy (anonymous):

so we can solve \[\frac{1}{3}=\frac{81-b^2}{9}\] for \(b\) i guess

OpenStudy (anonymous):

ok that was wrong, i mean solve \[\frac{1}{3}=\frac{\sqrt{81-b^2}}{9}\]

OpenStudy (anonymous):

i got ±9

OpenStudy (anonymous):

for b

OpenStudy (anonymous):

no i don't think so lets solve

OpenStudy (anonymous):

\[\frac{1}{3}=\frac{\sqrt{81-b^2}}{9}\] mulitply both sides by 9 to get \[3=\sqrt{81-b^2}\] square and get \[9=81-b^2\] so \[b^2=81-9=72\] making \(b^2=72\)

OpenStudy (anonymous):

final answer i think is \[\frac{x^2}{81}+\frac{y^2}{72}=1\] lets check it

OpenStudy (anonymous):

yes, it is right, here is the check note how i wrote "ellipse" first so it gives all the details http://www.wolframalpha.com/input/?i=ellipse+x^2%2F81%2By^2%2F72%3D1

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw !

OpenStudy (anonymous):

So i know how to get the center fine. but is it the same thing for the foci? to get a and b?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

the foci you get from \(c\)

OpenStudy (anonymous):

and \(c\) you get from \(a\) and \(b\) it is \(c^2=a^2-b^2\) or \(c=\sqrt{a^2-b^2}\)

OpenStudy (anonymous):

thats what i want to know, how to get a and b. thats always the part i get wrong

OpenStudy (anonymous):

really it depends on what you know in the first one we knew \(a=9\) and \[\frac{1}{3}=\frac{c}{a}=\frac{c}{9}\] we used that to find \(b^2\)

OpenStudy (anonymous):

oh ok, i see now. i thought they were different for ellipses and hyperbolas

OpenStudy (anonymous):

Find the standard form of the equation of the ellipse with the given characteristics. center: (-1, 7) a = 3c foci: (- 2, 7), (0, 7)\[\frac{(x+2)^2}{a^2}+\frac{(y-7)^2}{b^2}=1\]

OpenStudy (anonymous):

oops \[\frac{(x+1)^2}{a^2}+\frac{(y-7)^2}{b^2}=1\]

OpenStudy (anonymous):

now here the foci are 1 unit to the left and right of the center this tells you \(c=1\)

OpenStudy (anonymous):

and since you are told that \(a=3c\) given that \(c=1\) this gives you \(a=3\) and so \(a^2=9\)

OpenStudy (anonymous):

\[\frac{(x+2)^2}{9}+\frac{(y-7)^2}{b^2}=1\]

OpenStudy (anonymous):

now you need \(b^2\) and again \(a^2=b^2+c^2\) so \[9=b^2+1\] making \[b^2=8\]

OpenStudy (anonymous):

ok i think i got that, thanks!

OpenStudy (anonymous):

\[\frac{(x+1)^2}{9}+\frac{(y-7)^2}{8}=1\] http://www.wolframalpha.com/input/?i=ellipse+ \frac{%28x%2B1%29^2}{9}%2B\frac{%28y-7%29^2}{8}%3D1

OpenStudy (anonymous):

if you want to read a really good clear explanation of all of these, with good worked out examples, try this http://www.purplemath.com/modules/ellipse.htm

OpenStudy (anonymous):

ok thanks again

OpenStudy (anonymous):

or even this http://www.purplemath.com/modules/conics.htm yw